OpenStudy (anonymous):

Combine like terms. 5/6x - 2/3x A. 1/6x B.1/3x C.1/2x D. x

OpenStudy (anonymous):

@openstudyhelp98765432

OpenStudy (anonymous):

@DanJS

OpenStudy (xcecex):

Okay so, to make it more simple lets turn those fractions into decimals.

OpenStudy (xcecex):

5/6= 1.2 And 2/3= 1.5.

OpenStudy (xcecex):

Now we will subtract those.

OpenStudy (xcecex):

And 0.3= 1/3. And dont forget your variable so its 1/3x.

OpenStudy (anonymous):

Thanks you could you help me with one more @XceceX ?

OpenStudy (xcecex):

Of course! :)

OpenStudy (anonymous):

Thanks you ^-^

OpenStudy (xcecex):

Your welcome ^~^